# Deep learning-based species-area models reveal multi-scale patterns of species richness and turnover

### v1

@@ 5.1 Problem statement
$g_{\Theta}(m)=c+(S_{\text{tot}}-c)\left(1-\exp\left[-(m/e)^{b}\right]\right)$ (4)
where $c$ corresponds to the lower asymptote as $m\to 0$ (minimum expected species richness), $S_{\text{tot}}$ corresponds to the upper asymptote as $m\to\infty$ (total species richness), $b$ controls the shape parameter determining the rate of species accumulation, and $e$ is the scale parameter representing the characteristic sampling effort. As visually depicted in Fig. 1b, we found that this function provides a good fit for arbitrary rarefaction curves obtained with our dataset.
@@ 5.2 Deep SAR model
Parametrization of rarefaction curves.
Rarefaction curves are location and area-specific and depend on the environmental characteristics of the spatial unit considered (Dengler and Oldeland, 2010). To predict the species richness $S_{\text{obs}}$ of a spatial unit with arbitrary size and location, we condition $g_{\Theta}$ on features $X$ associated with the spatial unit, which include its area and statistics of environmental features within the spatial unit. We use a neural network $h_{\theta}$ to learn the mapping from the features $X$ to the parameters $\Theta$ of the rarefaction curve model $g_{\Theta}$:
@@
The predicted species richness for a given sampling effort $m$ is then
$\mathbb{E}[S_{\text{obs}}(m\mid X)]=g_{\Theta}(m,X)=g_{h_{\theta}(X)}(m).$ (6)
@@
$L(\theta)=\sum_{k=1}^{K}\left[g_{h_{\theta}(X_{k})}(m_{k})-S_{\text{obs}}^{(k)}(m_{k})\right]^{2},$ (7)
where $S_{\text{obs}}^{(k)}(m_{k})$ represents the observed species richness from $m_{k}$ randomly selected surveys within spatial unit $k$ with features $X_{k}$.
Training data generation.
We generated training samples by randomly placing spatial units containing at least two ecological surveys, with log-transformed areas following a uniform distribution between $10^{4}$ and $10^{11}$ m2. Within each spatial unit $k$ of area $A_{k}$, we identified the $M_{k}$ surveys from the EVA database whose coordinates fell within its bounds. We then selected a random subsample of $m_{k}$ ecological surveys, where $m_{k}$ followed a log-uniform discrete distribution $m_{k}\sim\mathcal{U}_{\log}\{2,3,...,M_{k}\}$. Based on these selected surveys, we calculated the observed species richness $S_{\text{obs}}^{(k)}$, which corresponded to the number of unique species observed across the selected ecological surveys. We assessed how the number of training samples influences model performance and report the results in Fig. S5. Model performance increased with the number of training samples, and we found that the optimal number is on the order of magnitude of the number of surveys used.
Model architecture and hyperparameters.
Our deep SAR models employed fully connected multi-layer perceptrons implemented in PyTorch. The network architecture consisted of 6 fully connected layers with 32, 128, 256, 256, 128, and 32 neurons respectively, using leaky ReLU activation functions. Input features included the spatial unit area and the mean and standard deviation of environmental variables within each spatial unit (detailed in Biodiversity data.). For ensemble predictions used in Contribution of area and environment across spatial scales and Scale-specific patterns of plant species richness and turnover in Europe, we trained 5 independent models and aggregated their outputs using arithmetic mean. Standard deviations of ensemble predictions are provided in Fig. S9. We used batch sizes of 1024 samples over 100 epochs for training, which we found sufficient for convergence as shown in Fig. S8. We used the Adam optimizer (Kingma and Ba, 2014) with decoupled weight decay regularization (Loshchilov and Hutter, 2019) and a learning rate scheduler with 20-epoch patience and 0.5 decay factor. Data was split into 90% training and 10% validation subsets for both ablation studies and final ensemble models.
@@ 5.3 Data
We used a dataset of 502,724 vegetation plots from the European Vegetation Archive (Chytrý et al. (2016); https://euroveg.org/eva-database/version/2023-02-04; project 172; data retrieved on 27 February 2023). We selected only plots with available area information and coordinate uncertainty less than 1 km. We additionally discarded plots not located on land within Europe and excluded non-vascular taxa from the dataset. Vegetation plots were analyzed without distinction of habitat types. The data filtering resulted in 488,622 vegetation plots for downstream analyses (see Fig. S10). Additionally, we used the Global Inventory of Floristic Taxonomy (GIFT) database (Weigelt et al., 2020), filtered to select records occurring within continental Europe, resulting in 184 exhaustive species records. We harmonized species names between the EVA and GIFT databases following procedures detailed at https://github.com/vboussange/DeepSAR. This resulted in the EVA-based dataset comprising 9,480 distinct species, while the GIFT-based dataset comprised 33,729 distinct species.
@@
We calculated environmental features using seven bioclimatic variables from the CHELSA-BIOCLIM+ dataset (Brun et al., 2022): mean annual air temperature (bio1), temperature seasonality (bio4), annual precipitation (bio12), precipitation seasonality (bio15), near-surface wind speed (sfcWind), potential evapotranspiration (pet), and surface downwelling shortwave radiation (rsds). These variables exhibit low pairwise correlations across the geographical extent (Spearman $\rho<0.5$). The associated mean climate features also show low correlations (Spearman $\rho<0.55$), whereas climate heterogeneity features display higher but still moderate correlations (Spearman $\rho<0.75$; see Fig. S7).
@@ 6 Data availability
The preprocessed EVA dataset with anonymized species names and the associated GIFT dataset anonymized with the same procedure are available in our GitHub repository at https://github.com/vboussange/DeepSAR. The original EVA dataset has restricted access and is available upon request at https://euroveg.org/eva-database. Scripts for downloading and processing the GIFT database are included in our GitHub repository. Upon manuscript acceptance, species richness rasters and trained model weights will be made publicly available.
@@ 8 Author contributions
V.B., P.B., J.T.M., G.M., J.P., T.S., N.E.Z., and D.N.K. conceived the study. V.B. developed the methodology, implemented the software, and performed the formal analysis. P.B., J.T.M., G.M., J.P., T.S., and D.N.K. contributed to methodological development. Data were provided by I.A., H.B., M.C., S.K., Z.L., M.V., I.B., K.T.E., J.L., and J.-C.S. V.B. wrote the original draft. All authors contributed to manuscript review and editing. N.E.Z. and D.N.K. acquired funding. D.N.K. supervised the project. Authors are listed alphabetically in the following order: the core team leading the study (V.B., P.B., J.T.M., G.M., J.P., T.S., N.E.Z.), followed by contributors from the FeedBaCks consortium (I.A., H.B., M.C., S.K., Z.L., M.V.) and EVA data custodians (I.B., K.T.E., J.L., J.-C.S.), also alphabetically within each group, and last the senior author.
@@ 9 Acknowledgements
Appendix S1 Supplementary materials
Derivation of the Michaelis-Menten model from the exponential distribution.
[…]
@@
Figure S9: Relative standard deviation of the predictions obtained from the deep SAR ensemble model and displayed in Fig. 4.

### v2

@@ 5.1 Estimating total species richness from small samples
$g_{\Theta}(a)=c+(S_{\mathrm{T}}-c)(1-\exp\left[-(a/e)^{b}\right]),$ (9)
which is at its minimum ($c$) at $a=0$ and converges to its maximum ($S_{\mathrm{T}}$) as $a$ becomes large. The offset $c$ captures the species richness at the smallest scale at the grain of the data. $b$ controls the shape parameter determining the rate of species accumulation, and $e$ is the scale parameter representing the characteristic sampling effort. As visually depicted in Fig.˜1b, we found that Eq.˜9 provides a good fit for arbitrary SARs obtained through our method and with the data we used.
@@ 5.2 Deep SAR model
Conditioning the rarefaction curve on environmental features and area of the spatial unit.
Rarefaction curves depend in general on the area and the environmental covariates associated with the spatial unit under consideration (Dengler and Oldeland, 2010). We make use of this dependence to map total species richness estimates over space. More precisely, to predict the observed species richness $S$ of a spatial unit $\mathsf{A}$} with arbitrary size and location, we condition $g_{\Theta}$ on features $X$ associated with the spatial unit, which include its area and statistics of environmental features within the spatial unit. We use a neural network $h_{\theta}$ to learn the mapping from the features $X$ to the parameters $\Theta$ of the rarefaction curve model $g_{\Theta}$:
@@
The predicted expected species richness for a given sampling effort $a$ is then
$\mathbb{E}[S(a)\mid X]=g_{\Theta}(a,X)=g_{h_{\theta}(X)}(a).$ (11)
@@
$L(\theta)=\sum_{k=1}^{K}\left[g_{h_{\theta}(X_{k})}(|\mathsf{a}_{{\boldsymbol{i}}_{k}}|)-S(\mathsf{a}_{{\boldsymbol{i}}_{k}})\right]^{2},\quad\mathsf{a}_{{\boldsymbol{i}}_{k}}\subset{\mathsf{A}}_{k},$ (12)
where $S(\mathsf{a}_{{\boldsymbol{i}}_{k}})$ represents the observed species richness from $m_{k}=|\mathsf{a}_{{\boldsymbol{i}}_{k}}|$ randomly selected surveys within spatial unit $k$ with features $X_{k}$.
Deep SAR model training.
We generated sample sets $\mathsf{a}_{{\boldsymbol{i}}_{k}}$ by randomly placing spatial units ${\mathsf{A}}_{k}$ containing at least two ecological surveys, with log-transformed area $\log|{\mathsf{A}}_{k}|$ following a uniform distribution between $10^{4}$ and $10^{11}$ m2. Within each spatial unit ${\mathsf{A}}_{k}$, we identified the $M_{k}$ surveys from the EVA database whose coordinates fall within its bounds. We then selected a random subsample of $m_{k}\leq M_{k}$ ecological surveys, where $m_{k}$ follows a log-uniform discrete distribution $m_{k}\sim\mathcal{U}_{\log}\{2,3,...,M_{k}\}$ and calculated their observed species richness. We assessed how the number of training samples influences model performance and report the results in Fig.˜S1. Model performance increases with the number of sample sets, and we found that the optimal number is on the order of magnitude of the number of surveys used.
Neural network architecture and hyperparameters.
Our deep SAR models employ fully connected multi-layer perceptrons implemented in PyTorch. The network architecture consists of 6 fully connected layers with 32, 128, 256, 256, 128, and 32 neurons respectively, using leaky ReLU activation functions. Input features include the spatial unit area and the mean and standard deviation of environmental variables within each spatial unit (detailed in Biodiversity data.). For ensemble predictions used in Contribution of area and environment across spatial scales and Scale-specific patterns of plant species richness and turnover in Europe, we trained 5 independent models and aggregated their outputs using arithmetic mean. Standard deviations of ensemble predictions are provided in Fig.˜S6. We used batch sizes of 1024 samples over 100 epochs for training, which we found sufficient for convergence as shown in Fig.˜S5. We used the Adam optimizer (Kingma and Ba, 2014) with decoupled weight decay regularization (Loshchilov and Hutter, 2019) and a learning rate scheduler with 20-epoch patience and 0.5 decay factor. Data was split into 90% training and 10% validation subsets for both ablation studies and final ensemble models.
@@ 5.3 Data
We used a dataset of 502,724 vegetation plots from the European Vegetation Archive (Chytrý et al. (2016); https://euroveg.org/eva-database/version/2023-02-04; project 172; data retrieved on 27 February 2023). We selected only plots with available area information and coordinate uncertainty less than 1 km. We additionally discarded plots not located on land within Europe and excluded non-vascular taxa from the dataset. Vegetation plots were analyzed without distinction of habitat types. The data filtering resulted in 488,622 vegetation plots for downstream analyses (see Fig.˜S7). Additionally, we used the Global Inventory of Floristic Taxonomy (GIFT) database (Weigelt et al., 2020), filtered to select surveys occurring within continental Europe, resulting in 184 exhaustive species surveys. We harmonized species names between the EVA and GIFT databases following procedures detailed at https://github.com/vboussange/DeepSAR. This resulted in the EVA-based dataset comprising 9,480 distinct species, while the GIFT-based dataset comprised 33,729 distinct species.
@@
We calculated environmental features using seven bioclimatic variables from the CHELSA dataset (Karger et al., 2021; Brun et al., 2022): mean annual air temperature (bio1), temperature seasonality (bio4), annual precipitation (bio12), precipitation seasonality (bio15), near-surface wind speed (sfcWind), potential evapotranspiration (pet), and surface downwelling shortwave radiation (rsds). These variables exhibit low pairwise correlations across the geographical extent (Spearman $\rho<0.5$). The associated mean climate features also show low correlations (Spearman $\rho<0.55$), whereas climate heterogeneity features display higher but still moderate correlations (Spearman $\rho<0.75$; see Fig.˜S3).
@@ 6 Data availability
The preprocessed EVA dataset with anonymized species names, the associated GIFT dataset anonymized with the same procedure, the pretrained weights of the deep SAR ensemble model, and tutorials for how to use it, are available under our GitHub repository at https://github.com/vboussange/DeepSAR. The original EVA dataset has restricted access and is available upon request at https://euroveg.org/eva-database. Climate data used in this study is available from the CHELSA database https://chelsa-climate.org. Scripts for downloading and processing the original GIFT database are included in our GitHub repository.
@@ 8 Author contributions
V.B., B.W., P.B., J.T.M., G.M., J.P., T.S., N.E.Z., and D.N.K. conceived the study. V.B. developed the methodology and implemented the software. V.B and B.W. performed the formal analysis. P.B., J.T.M., G.M., J.P., T.S., and D.N.K. contributed to methodological development. Data were provided by I.A., H.B., M.C., S.K., Z.L., M.V., I.B., K.T.E., J.L., and J.-C.S. V.B. wrote the original draft. All authors contributed to manuscript review and editing. N.E.Z. and D.N.K. acquired funding. D.N.K. supervised the project. Authors are listed alphabetically in the following order: the core team leading the study (V.B., P.B., J.T.M., G.M., J.P., T.S., N.E.Z.), followed by contributors from the FeedBaCks consortium (I.A., H.B., M.C., S.K., Z.L., M.V.) and EVA data custodians (I.B., K.T.E., J.L., J.-C.S.), also alphabetically within each group, and last the senior author (D.N.K.).
@@ 9 Acknowledgements
Appendix S1 Supplementary text
S1.1 Decomposition of $p_{0}$
Step 1. Definition.
$p_{0}(a):=\mathbb{P}\!\left(N(a)=0\,\mid\,N({\mathsf{A}})\geq 1\right).$ (S1)
Step 2. Law of total probability for conditional probabilities.
$\mathbb{P}\!\left(N(a)=0\,\mid\,N({\mathsf{A}})\geq 1\right)=\sum_{n\geq 1}\mathbb{P}\!\left(N(a)=0\mid N({\mathsf{A}})=n\right)\;\mathbb{P}\!\left(N({\mathsf{A}})=n\mid N({\mathsf{A}})\geq 1\right).$ (S2)
Step 3. Identify ZT–SAD and spatial factor.
$f_{n}^{+}\;:=\;\mathbb{P}\!\left(N({\mathsf{A}})=n\mid N({\mathsf{A}})\geq 1\right),\qquad q_{n}(a)\;:=\;\mathbb{P}\!\left(N(a)=0\mid N({\mathsf{A}})=n\right),$ (S3)
so
$p_{0}(a)\;=\;\sum_{n\geq 1}f_{n}^{+}q_{n}(a).$ (S4)
Step 4. Special case: uniform random placement.
If $q_{n}(a)=(1-\alpha)^{n}$ with $\alpha=a/|{\mathsf{A}}|$ and $G\{f^{+}_{n}\}(z)=\sum_{n\geq 1}f_{n}^{+}z^{n},$ then
$p_{0}(a)=G\{f^{+}_{n}\}(1-\alpha),$ (S5)
which matches the result of Coleman (1981).
Appendix S2 Supplementary table
SAD (ZT) $\boldsymbol{f^{+}_{n}}$ (PMF over species) $\boldsymbol{G\{f^{+}_{n}\}(z)}$ (PGF of SAD) $\boldsymbol{\mathcal{S}(\alpha)}$ Refs
$\mathrm{Logseries}(x)$ $f^{+}_{n}=\dfrac{x^{n}}{n\,[-\ln(1-x)]},\;n\geq 1$ $G\{f^{+}_{n}\}(z)=\dfrac{\ln(1-xz)}{\ln(1-x)}$ $\mathcal{S}(\alpha)=S_{\mathrm{T}}\!\bigl{[}1-\dfrac{\ln\!\bigl{(}1-x(1-\alpha)\bigr{)}}{\ln(1-x)}\bigr{]}$ Coleman (1981); He and Legendre (2002)
$\mathrm{Pois}^{+}(\mu)$ $f^{+}_{n}=\dfrac{e^{-\mu}\mu^{n}/n!}{1-e^{-\mu}},\;n\geq 1$ $G\{f^{+}_{n}\}(z)=\dfrac{e^{\mu(z-1)}-e^{-\mu}}{1-e^{-\mu}}$ $\mathcal{S}(\alpha)=S_{\mathrm{T}}\,\dfrac{1-e^{-\mu\alpha}}{1-e^{-\mu}}$
$\mathrm{NB}^{+}(r,p)$ $f^{+}_{n}=\dfrac{\binom{n+r-1}{n}(1-p)^{n}p^{r}}{1-p^{r}},\;n\geq 1$ $G\{f^{+}_{n}\}(z)=\dfrac{\left(\dfrac{p}{1-(1-p)z}\right)^{r}-p^{r}}{1-p^{r}}$ $\mathcal{S}(\alpha)=S_{\mathrm{T}}\,\dfrac{1-\left(\dfrac{p}{\,p+(1-p)\alpha\,}\right)^{r}}{1-p^{r}}$ He and Legendre (2002)
$\mathrm{Geom}^{+}(p)^{*}$ $f^{+}_{n}=p(1-p)^{\,n-1},\;n\geq 1$ $G\{f^{+}_{n}\}(z)=\dfrac{p\,z}{1-(1-p)z}$ $\mathcal{S}(\alpha)=S_{\mathrm{T}}\,\dfrac{\alpha}{\,p+(1-p)\alpha\,}$ He and Legendre (2002)
Most even $f^{+}_{n^{\ast}}=1\ \text{with }n^{\ast}=N/S_{\mathrm{T}}\in\mathbb{N};\ \text{rest }0\$$G\{f^{+}_{n}\}(z)=z^{\,N/S_{\mathrm{T}}}$ $\mathcal{S}(\alpha)=S_{\mathrm{T}}\bigl{[}1-(1-\alpha)^{N/S_{\mathrm{T}}}\bigr{]}$ He and Legendre (2002)
Most uneven $f^{+}_{1}=\dfrac{S_{\mathrm{T}}-1}{S_{\mathrm{T}}},\ f^{+}_{\,N-S_{\mathrm{T}}+1}=\dfrac{1}{S_{\mathrm{T}}};\ \text{rest }0\$$G\{f^{+}_{n}\}(z)=\dfrac{S_{\mathrm{T}}-1}{S_{\mathrm{T}}}\,z+\dfrac{1}{S_{\mathrm{T}}}\,z^{\,N-S_{\mathrm{T}}+1}$ $\mathcal{S}(\alpha)=1+(S_{\mathrm{T}}-1)\alpha-(1-\alpha)^{\,N-S_{\mathrm{T}}+1}$ He and Legendre (2002)
Table S1: Expected species richness for uniform random placement for a selection of SADs, with the event of an individual falling into $a$ being a Bernoulli trial with probability parameter $\alpha:=a/|A|$. All formulas use $\mathcal{S}(\alpha)=\mathbb{E}[S(\alpha)]=S_{\mathrm{T}}\bigl{[}1-G\{f^{+}_{n}\}(1-\alpha)\bigr{]}$. SAD = species abundance distribution; ZT = zero-truncated. Note that $\mathrm{Geom}^{+}(p)=\mathrm{NB}^{+}(1,p)$.
@@
Figure S4: Relative bias of the deep SAR model on the EVA test dataset, calculated as the ratio of the difference between predicted and observed species richness to the observed species richness.
@@
Figure S6: Relative standard deviation of the predictions obtained from the deep SAR ensemble model and displayed in Fig.˜4.